OpenStudy (anonymous):

At what interest rate the must $4500 be compounded annually to equal $10,174.07 after 14 years? Thanks! ;)

OpenStudy (anonymous):

use V = P (1 + r)^y 10,174.07 is v 14 is y 4500 is p plug these in to find r 10174.07=4500(1+r)^14

OpenStudy (anonymous):

@CalebBeavers Hi, can you help me to find r I am stuck... Thanks!!

OpenStudy (anonymous):

it will be 2.26

OpenStudy (anonymous):

10174.07/4500=(1+r)^14 \[\large \sqrt[14]{10174.07/4500}-1=r\]

OpenStudy (anonymous):

Not sure what I am doing wrong...

OpenStudy (anonymous):

First you divide both sides by 4500, then take the 14th root of both sides, and then subtract 1 from both sides

OpenStudy (anonymous):

Ok! let me try..

OpenStudy (anonymous):

How can I plug the 14th square root on my TI-30xIIS calculator, Do you have an idea? Thanks!

OpenStudy (anonymous):

Not sure...there might be a button that looks like \[\sqrt[y]{x}\large\]

OpenStudy (anonymous):

Yes, there is one like this \[\sqrt[x]{?}\]

OpenStudy (anonymous):

Its there is an easy way to do this??